OpenStudy (anonymous):

I'm trying to understand how to reduce Gamma functions for Laplace transforms, starting from Gamma(n/2)=sqrt(pi/2). More detail with formatting below:

OpenStudy (anonymous):

\[\Gamma (\frac{5}{2})=(\frac{3}{2}) \Gamma (\frac{3}{2}) =(\frac{1}{2}) ( \frac{3}{2})\Gamma (\frac{1}{2}) =\frac{3 }{4} \sqrt{\pi }\]

OpenStudy (anonymous):

It's the reduction of the formula above that's I'm not understanding.

OpenStudy (anonymous):

Ah, I think I have it. Is it because of this? (See image).

OpenStudy (unklerhaukus):

\[\Gamma(z+1) = z\cdot \Gamma (z)\]

OpenStudy (anonymous):

Which gives me this, right?

OpenStudy (unklerhaukus):

that looks right, [ i haven't practiced these in a while ]

OpenStudy (unklerhaukus):

for reducing the Gamma function, this form is used: \[\Gamma(z)=(z-1)\Gamma(z-1)\]

OpenStudy (anonymous):

Ah, that'll be handy. Thank you, I'll try it on this next problem.

OpenStudy (unklerhaukus):

\[\sqrt\pi=\Gamma(1/2)=2\cdot \Gamma(3/2)=2\times\frac23\cdot \Gamma(5/2)=2\times \frac23\times\frac25\cdot\Gamma(7/2)=\cdots \]

OpenStudy (anonymous):

Excellent. We didn't do a lot of this in class, but I didn't want to be blindsided on the final tomorrow, either.

OpenStudy (anonymous):

Thanks again.

OpenStudy (unklerhaukus):

ps wolfram agrees with your Laplace transform https://www.wolframalpha.com/input/?i=Laplace+transform+%5Bt%5E%281%2F2%29%2B3t%5D

OpenStudy (anonymous):

I love it when Wolfram agrees. :)

OpenStudy (unklerhaukus):

if you can remember only one of the forms of the gamma function, make sure to manipulate it to be in a useful form before you try to use it, because if you're trying to use is in backwards it gets confusing

OpenStudy (anonymous):

Cool. Conveniently, my calculator can do them pretty easily. It should be pretty easy to check my work.